Shawn and his bike have a total mass of
57.9 kg. Shawn rides his bike 0.91 km in
12.5 min at a constant velocity.
The acceleration of gravity is 9.8 m/s2.
What is Shawn’s kinetic energy?
Answer in units of J.
Don't round answer
1
Expert's answer
2019-11-04T09:54:54-0500
"v=\frac{s}{t}=\frac{910}{12.5\cdot 60}=1.213\frac{m}{s}"

Shawn’s kinetic energy:


"K=0.5mv^2=0.5(57.9)(1.213)^2=42.6\ J"